OpenStudy (anonymous):

HELP PLEASE!! The point Three point five comma two point five would not fall on the line for which of the following representations? Nine minus two y equals two x minus three. four y equals negative seven x minus three point five. Five minus four x equals negative two y minus four. Negative three y equals negative two x minus zero point five.

OpenStudy (anonymous):

(3.5,2.5) A) 9 - 2y = 2x - 3 B) 4y = -7x - 3.5 C) 5 - 4x = -2y - 4 D) -3y = -2x - 0.5 Answer is C for, 5 -4x = -9 and -2y - 4 = -9 -9 = -9

OpenStudy (anonymous):

so A. 9- 2y + 2x - 3 B. 4y + -7x - 3.5 C. 5 - 4x + 2y- 4 D. -3y = -2x - 0.5 if i were you i would choose C.